OpenStudy (anonymous):

Help please... Let f(x)=-2x+4 and g(x)=-6x-7. find f(x)-g(x)

OpenStudy (anonymous):

put them in parentheses, then put a minus sign between them

OpenStudy (anonymous):

(-2x+4)-(-6x-7) like this? @satellite73

OpenStudy (anonymous):

yes then distribute the minus sign for the second parentheses, and combine like terms

OpenStudy (anonymous):

4x-3

OpenStudy (anonymous):

close but you did not distribute completely

OpenStudy (anonymous):

\[(-2x+4)-(-6x-7) =-2x+4+6x+7\]

OpenStudy (anonymous):

4x+11 I am confused

OpenStudy (anonymous):

your second answer is correct

OpenStudy (anonymous):

when you distribute the minus sign, you have to change the sign if each term you see

OpenStudy (anonymous):

Oh okay thank you! can you help me with 2 more questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

heres one what is the inverse of the given relation? y=3x+12

OpenStudy (anonymous):

wow we really stepped it up from the last one!

OpenStudy (anonymous):

can you solve \[y=3x+12\] for \(x\)?

OpenStudy (anonymous):

Yeah. 12+y=3x 4y=x how would i get y alone now.

OpenStudy (anonymous):

hmm lets start at the beginning

OpenStudy (anonymous):

I know i have to get x alone

OpenStudy (anonymous):

\[y=3x+12\] and we want to solve for \(x\) before i do it, let me ask you a question do you have answer choices? or do you have to come up with it on your own i just want to know what the choices look like before we start

OpenStudy (anonymous):

i have to type it all out i dont have answer choices.

OpenStudy (anonymous):

ok i was asking because i don't know if your answer is suppose to look like \(y=something\) or \(x=something\) but not matter lets solve for \(x\)

OpenStudy (anonymous):

\[y=3x+12\] solve for \(x\) in two steps 1) subtract \(12\) from both sides 2) divide both sides by \(3\)

OpenStudy (anonymous):

\[y=3x+12\\ y-12=3x\\ \frac{y-12}{3}=x\]

OpenStudy (anonymous):

i know you get 4

OpenStudy (anonymous):

i dont know where to put it lol

OpenStudy (anonymous):

the reason i was asking the previous question is because i dont know if your answer is supposed bo be \[y=\frac{x-12}{3}\] or \[x=\frac{y-12}{3}\]

OpenStudy (anonymous):

i am not sure what you mean

OpenStudy (anonymous):

i was dividing but now i get it.

OpenStudy (anonymous):

yeah when i say "divide' i don't really mean do a division, i just mean write it

OpenStudy (anonymous):

Oh okay. I have one more i need help on. rewrite 3^sqrt(27x-81-5) to make it easy to graph using a translation.

OpenStudy (anonymous):

is this really \[\sqrt[3]{27x^2-81-5}\]? seen unlikely

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i have no idea repost this as a new question and maybe you will get a good answer

OpenStudy (anonymous):

I figured it out thank you though!